Three charged particles are placed at the corners of an equilateral triangle of side 1.20 m. The charges are Q1= 6.7 μC, Q2 = -9.4 μC, Q3= -4.9 μC .


  1. Calculate the magnitude of the net force on particle 2 due to the other two.
  2. Calculate the magnitude of the net force on particle 3 due to the other two.
  3. Calculate the direction of the net force on particle 3 due to the other two.
1
Expert's answer
2021-02-12T16:17:27-0500

1) Let charge "Q_1" placed at the top vertice of the equilateral triangle and charges "Q_2" and "Q_3" placed in the corners of the base of the equilateral triangle. Let the side "Q_1Q_2" be the "x"-axis. The net force on charge "Q_2" is the vector sum of forces "F_{12}" and "F_{32}" due to charges "Q_1" and "Q_3".

Let’s find the magnitudes of these forces:


"F_{12}=\dfrac{k|Q_1Q_2|}{r^2},""F_{12}=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot|6.7\cdot10^{-6}\ C\cdot(-9.4\cdot10^{-6}\ C)|}{(1.2\ m)^2}=0.394\ N,""F_{32}=\dfrac{k|Q_3Q_2|}{r^2},""F_{32}=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot|(-4.9\cdot10^{-6}\ C)\cdot(-9.4\cdot10^{-6}\ C)|}{(1.2\ m)^2}=0.288\ N."

Then, we can find the projections of forces "F_{12}" and "F_{32}" on axis "x" and "y":


"F_x=F_{12x}+F_{32x}=-0.394\ N-0.288\ N\cdot cos60^{\circ}=-0.538 N,""F_y=F_{12y}+F_{32y}=0-0.288\ N\cdot sin60^{\circ}=-0.249 N."

Finally, the net electric force can be found from the Pythagorean theorem:


"F_{net}=\sqrt{F_x^2+F_y^2},""F_{net}=\sqrt{(-0.538\ N)^2+(-0.249\ N)^2}=0.593\ N."

2) Let the side "Q_1Q_3" be the "x"-axis. The net force on charge "Q_3" is the vector sum of forces "F_{13}" and "F_{23}" due to charges "Q_1" and "Q_2".

Let’s find the magnitudes of these forces:


"F_{13}=\dfrac{k|Q_1Q_3|}{r^2},""F_{13}=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot|6.7\cdot10^{-6}\ C\cdot(-4.9\cdot10^{-6}\ C)|}{(1.2\ m)^2}=0.205\ N,""F_{23}=\dfrac{k|Q_2Q_3|}{r^2},""F_{23}=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot|(-9.4\cdot10^{-6}\ C)\cdot(-4.9\cdot10^{-6}\ C)|}{(1.2\ m)^2}=0.288\ N."

Then, we can find the projections of forces "F_{13}" and "F_{23}" on axis "x" and "y":


"F_x=F_{13x}+F_{23x}=-0.205\ N+0.288\ N\cdot cos60^{\circ}=-0.061 N,""F_y=F_{13y}+F_{23y}=0-0.288\ N\cdot sin60^{\circ}=-0.249 N."

Finally, the net electric force can be found from the Pythagorean theorem:


"F_{net}=\sqrt{F_x^2+F_y^2},""F_{net}=\sqrt{(-0.061\ N)^2+(-0.249\ N)^2}=0.256\ N."

3) We can find the direction of the net force on charge "Q_3" as follows:


"cos\theta=\dfrac{F_x}{F_{net}},""\theta=cos^{-1}(\dfrac{F_x}{F_{net}})=cos^{-1}(\dfrac{-0.061 N}{0.256\ N})=104^{\circ}."